OpenStudy (anonymous):

Find the value of x if the area of the triangle is 297.5 in2?

OpenStudy (anonymous):



OpenStudy (anonymous):

the base is 70 inches in case you can read it

OpenStudy (anonymous):

cant*

OpenStudy (anonymous):

A= b*h/2

OpenStudy (anonymous):

so what reverse it?

OpenStudy (anonymous):

297.5 = 70*h/2 now solve for h

OpenStudy (anonymous):

297.5=70*h/2 -70 -70 227.5=h/2 /2 /2 113.75=h ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

(297.5) *2 =70 *h

OpenStudy (anonymous):

h= (295)*2/70

OpenStudy (anonymous):

im confused ha

OpenStudy (anonymous):

what do i use to solve for x

OpenStudy (anonymous):

sorry x=h it is the heigh

OpenStudy (anonymous):

ok thanks!